OpenStudy (anonymous):

Given the expression 3x2y + 4xy + 1 – 5x4y – 6y, do the following as instructed below: Write the polynomial in descending order. Classify the polynomial by the number of terms. State the degree of the polynomial.

OpenStudy (danjs):

every variable is to the first power?

OpenStudy (anonymous):

i dont get it

OpenStudy (danjs):

you can combine things with the same variable terms.. like 3x2y + 4xy - 5x4y = 6xy + 4xy - 20 xy = (6 + 4 - 20)xy = -10xy

OpenStudy (anonymous):

does it apply for a.. b. nd c

OpenStudy (danjs):

yes

OpenStudy (danjs):

need to simplify it first by combining the like terms

OpenStudy (danjs):

oh you know what, are those powers? like the first term \[3x^2*y\] ?

OpenStudy (anonymous):

NOPE . i dont know this at allllll

OpenStudy (danjs):

They probably are, usually put a x^2 hat for powers

OpenStudy (danjs):

So the prob is... \[3x^2y+4xy + 1 -5x^4 y - 6y\]

OpenStudy (danjs):

Descending order - Largest powers of x first, Order the terms in decreasing powers of x

OpenStudy (danjs):

Number of terms, are how many things are added or subtracted together

OpenStudy (danjs):

the degree is the highest power

OpenStudy (anonymous):

so 5x4y is the degree

OpenStudy (anonymous):

what does it mean by classify for B

OpenStudy (danjs):

sorry, the degree for terms like x^4*y is the sum of both variables exponents, here 4 + 1 = 5, the degree is 5

OpenStudy (danjs):

just means how many terms are in the polynomial, here 5

OpenStudy (anonymous):

what is the polynomial by the number of terms

OpenStudy (danjs):

each term is added or subtracted together

OpenStudy (danjs):

like 3x^2*y is one term

OpenStudy (princesssleelee):

@DanJS can you explain this to me please? i have the same question